OpenStudy (anonymous):

what is the value of x in the following equation? 4(x + 3) = 9x + 16 + 3x

OpenStudy (anonymous):

x= -1/2 or -0.5

OpenStudy (anonymous):

-1/2 is an option

OpenStudy (anonymous):

4x +12 = 9x+ 16+3x 4x - 9x - 3x = 16-12 -8x = 4 x = -1/2

OpenStudy (anonymous):

whats the value of x in the following equation? (x + 4) + 3 = 2x( + )